# Generalized Carlos Scales

### v4

@@ 2. Deriving the original Carlos scales and some new Carlos-type scales
Remark 1.
One anonymous referee pointed out that perhaps the simplest way of obtaining the alpha, beta, and gamma scales is by dividing the just perfect fifth (3:2 ratio) into equal parts: 9 equal parts for alpha, 11 equal parts for beta, and 20 equal parts for gamma. This implies the alpha unit is 77.995 cents, the beta unit is 63.814 cents, and the gamma unit is 35.09775 cents. These values differ ever so slightly from the alpha, beta, and gamma units as derived by Benson’s method (77.965, 63.833, and 35.099 cents respectively). To the human ear, these discrepancies would be imperceptible, but is nonetheless of potential mathematical interest. By dividing the just perfect fifth, the Carlos scales would, by definition, have a perfectly in tune fifth, while the Benson method yields scales with ever so slightly wide or narrow fifths. Carlos [4, p. 2] hints that she intended the perfect fifth to be exactly just, i.e. an exact $3:2$ ratio, by writing, “If you try to play through a one octave scale of Alpha, you’d find there are 4 steps to the minor third, 5 steps to the major third and 9 steps to the perfect (no kidding) fifth, …” Elsewhere in Carlos’s writings [3, 4], she uses the values 78.0, 63.8, and 35.1 cents respectively for unit size in the alpha, beta, and gamma scales: thus to three significant figures, there is no difference between the equal division of the just perfect fifth and Benson’s method. We proceed using Benson’s method.
@@
Next, one needs to experiment with various values of $a$ and $b$, and check to see if all the approximations are within an acceptable tolerance, say, $5$ cents. The author programmed Mathematica to perform such a search. Recall that the just minor third, major third, and perfect fifth are $315.64$, $386.31$, and $701.96$ cents, respectively. Thus, if
@@
Another related scale noted by Carlos [3] is the $\alpha^{\prime}$ scale, which in our notation is the $(8,10)$-Carlos scale. It is simply the alpha scale with each unit cut in half (or equivalently, the number of units is doubled). Here the unit is about $38.9825$ cents, half that of the alpha scale. Of course, the thirds and the fifth are identical to that of the alpha scale; but this “doubled alpha” scale has one advantage over the alpha scale: an octave requires about 30.783 units. Thus the octave approximated by the 31st note is about 8.45734 cents wide, which is much better than the 15th note of the alpha scale. Nonetheless, if $\gcd(a,b)>1$, that $(a,b)$-Carlos scale could reasonably be considered a trivial variation of the $\big{(}a/\gcd(a,b),b/\gcd(a,b)\big{)}$-Carlos scale, and therefore not of genuine interest. [^2] As usual, we denote the greatest common divisor of $a$ and $b$ by $\gcd(a,b)$.
@@
Similarly of particular note for an excellent octave in addition to the thirds and the fifth is the $(17,21)$-Carlos scale, as 65 units is 0.520 cents wider than an octave. The $(19,23)$-Carlos scale in addition to excellent thirds and fifths also has a very pleasing octave: 72 units is 3.56 cents wider than a perfect octave. Please see the table below for additional examples of $(a,b)$-Carlos scales of interest, all with less than 81 units per octave.
@@
3. Further Generalizations with four or more parameters or with alternative intervals optimized
@@ 3.1. Three or more independent parameters
Remark 2.
One of the anonymous referees pointed out that it would be worthwhile to acknowledge here the shortcomings of scales where the approximation to $6/5$ plus the approximation to $5/4$ does not exactly equal the interval approximating $3/2$; and that this phenomenon is called inconsistency, “and can make traditional composing techniques with chords frightfully awkward.”
@@ 3.2. Alternative intervals optimized
Note that unit size $x_{\{\mathrm{P4,P5}\}}^{(5,7)}=(15+2\log_{2}3)/218\approx 0.0833483\approx 0.08333333\dots=1/12$ gives a very close approximation to 12-ET.
Remark 3.
One could argue, however, that the preceding example is an abuse of Carlos’ original idea, as it could be considered wasteful to optimize for both the perfect fourth and fifth, as each is the inversion of the other. Indeed, the most simple solution $(a,b)=(5,7)$ does not really get us anything new, but rather a slightly out of tune version of 12-ET(!)
@@ Acknowledgments
The author wishes to thank Robert Schneider for his assistance and encouragement, and for valuable suggestions. The author also thanks the Mathematics and Music Lab at Michigan Technological University for invaluable consultation work. Finally, the author thanks the anonymous referees for catching a typographical error in the original submission, and for making many useful suggestions that improved the paper considerably.

### v5

@@ 2. Deriving the original Carlos scales and some new Carlos-type scales
We note that perhaps the simplest way of obtaining the alpha, beta, and gamma scales is by dividing the just perfect fifth (3:2 ratio) into equal parts: 9 equal parts for alpha, 11 equal parts for beta, and 20 equal parts for gamma. This implies the alpha unit is 77.995 cents, the beta unit is 63.814 cents, and the gamma unit is 35.09775 cents. These values differ ever so slightly from the alpha, beta, and gamma units as derived by Benson’s method (77.965, 63.833, and 35.099 cents respectively). To the human ear, these discrepancies would be imperceptible, but is nonetheless of potential mathematical interest. By dividing the just perfect fifth, the Carlos scales would, by definition, have a perfectly in tune fifth, while the Benson method yields scales with ever so slightly wide or narrow fifths. Carlos [4, p. 2] hints that she intended the perfect fifth to be exactly just, i.e. an exact $3:2$ ratio, by writing, “If you try to play through a one octave scale of Alpha, you’d find there are 4 steps to the minor third, 5 steps to the major third and 9 steps to the perfect (no kidding) fifth, …” Elsewhere in Carlos’s writings [3, 4], she uses the values 78.0, 63.8, and 35.1 cents respectively for unit size in the alpha, beta, and gamma scales: thus to three significant figures, there is no difference between the equal division of the just perfect fifth and Benson’s method. We proceed using Benson’s method.
@@
Next, one needs to experiment with various values of $a$ and $b$, and check to see if all the approximations are within an acceptable tolerance $v$, say, $v=$ 55 cents. While $5$ cents may be considered an arbitrary cutoff and that additional scales of interest would emerge if we increased the tolerance to $8$ cents, it seems that the conventional wisdom is that the trained human ear is sensitive to pitch differences of greater than or equal to $5$ cents; see, e.g. [6, p. 287, first full paragraph], so this seems to be a reasonable, though perhaps arbitrary cutoff to employ for our present study. Nonetheless, if we were to increase the limit to $5$, the fifths remain surprisingly stable and always around $702$ cents. The thirds fluctuate a little more. Remember that the distance from the perfect major third $(5/4)$ is $14$ cents for the tempered major third $(2^{1/3})$ and $22$ cents for the Pythagorean major third $(81/64)$. And the distance from the perfect minor third $(6/5)$ is about $16$ cents for the minor third of the meantone temperament $(4/5^{3/4})$ or for the tempered minor third and $22$ cents for the Pythagorean third $(32/27)$. In musical reality, the gaps therefore seem much greater than in the theory of the article. If we go, for example, from $v<5$ to $v<8$, we see that new scales appear, e.g., (11,13)-Carlos close to 41-ET, (15,19)-Carlos close to 48-ET.
The author programmed Mathematica to perform such a search. Recall that the just minor third, major third, and perfect fifth are $315.64$, $386.31$, and $701.96$ cents, respectively. Thus, if
@@
Another related scale noted by Carlos [3] is the $\alpha^{\prime}$ scale, which in our notation is the $(8,10)$-Carlos scale. It is simply the alpha scale with each unit cut in half (or equivalently, the number of units is doubled). Here the unit is about $38.9825$ cents, half that of the alpha scale. Of course, the thirds and the fifth are identical to that of the alpha scale; but this “doubled alpha” scale has one advantage over the alpha scale: an octave requires about 30.783 units. Thus the octave approximated by the 31st note is about 8.45734 cents wide, which is much better than the 15th note of the alpha scale. Nonetheless, if $\gcd(a,b)>1$, that $(a,b)$-Carlos scale could reasonably be considered a trivial variation of the $\big(a/\gcd(a,b),b/\gcd(a,b)\big)$-Carlos scale, and therefore not of genuine (mathematical) interest. [^2] As usual, we denote the greatest common divisor of $a$ and $b$ by $\gcd(a,b)$.
Nonetheless, there is indeed musical interest to consider $(a,b)$-Carlos scales where $\gcd(a,b)>1$. For example, the $(8,10)$-Carlos scale, called $\alpha^{\prime}$ by Carlos, is close to 31-ET, the $(10,12)$-Carlos scale ($\beta^{\prime}$) is close to 38-ET, $(12,15)$-Carlos ($\alpha^{\prime\prime}$) is close to 48-ET, etc. Further, one could ask the opposite question, namely for a given $N$-ET temperament, which $a$ and $b$ give the closet $(a,b)$-Carlos approximation? For example, the Hölder temperament (53-ET) is well approximated by $(14,17)$-Carlos; see further comments on this below.
@@
Similarly of particular note for an excellent octave in addition to the thirds and the fifth is the $(17,21)$-Carlos scale, as 65 units is 0.520 cents wider than an octave. The $(19,23)$-Carlos scale in addition to excellent thirds and fifths also has a very pleasing octave: 72 units is 3.56 cents wider than a perfect octave. Please see the table below for additional examples of $(a,b)$-Carlos scales of interest, all with less than 81 units per octave. The author does not intend to imply that there are no interesting examples with greater than 81 units per octave; the table just contains examples to suggest possibilities. For example, the $(25,31)$-Carlos scale is close to the sixteenth tone space (96-ET) often used in microtonal music.
@@
3. Further Generalizations with three or more parameters or with alternative intervals optimized
@@ 3.1. Three or more independent parameters
It is worth acknowledging here the shortcomings of scales where the approximation to $6/5$ plus the approximation to $5/4$ does not exactly equal the interval approximating $3/2$; and that this phenomenon is called inconsistency, and can make traditional composing techniques with chords awkward.
@@ 3.2. Alternative intervals optimized
Note that unit size $x_{\{\mathrm{P4,P5}\}}^{(5,7)}=(15+2\log_{2}3)/218\approx 0.0833483\approx 0.08333333\dots=1/12$ gives a very close approximation to 12-ET. One could argue, however, that the preceding example is an abuse of Carlos’ original idea, as it could be considered wasteful to optimize for both the perfect fourth and fifth, as each is the inversion of the other. Indeed, the most simple solution $(a,b)=(5,7)$ does not really get us anything new, but rather a slightly out of tune version of 12-ET(!)
@@ Acknowledgments
The author wishes to thank Robert Schneider and Eden Sills for their assistance and encouragement, and for valuable suggestions. The author also thanks the Mathematics and Music Lab at Michigan Technological University for invaluable consultation work. Finally, the author thanks the anonymous referees for catching a typographical error in the original submission, and for making many useful suggestions that improved the paper considerably.